OpenStudy (anonymous):

* Give an example of the commutative property of addition. So lost....don't remember this :/

OpenStudy (anonymous):

a + b = b + a

OpenStudy (anonymous):

in numbers: 4+9=13=9+4

OpenStudy (anonymous):

hey email me when u can ivy!